# Stagewise Reinforcement Learning and the Geometry of the Regret Landscape

### v1

@@ F.1 Asymptotically Unbiased Estimators
This follows from Theorem E.43. Indeed, by the weak law of large numbers
@@
where the expectation value is taken with respect to the generalized tempered posterior to $U$: $y_{j}\sim\Omega_{n,\beta}|_{U}$. Theorem E.43 then tells us that this expected value converges to $\lambda(U)$ in probability as $n\to\infty$, as required. ∎
F.2 The Annealed Posterior
Let’s discuss a variant of the WBIC-based LLC estimator with the same limiting properties as we saw in Proposition F.4. We will replace the tempered posterior in our definition with the annealed posterior (see Seung et al. [1992] for a discussion of this concept from its origin in statistical mechanics).
Definition F.5.
@@ F.2 The Annealed Posterior
Proposition F.7.
Fix a log resolution and a chart $U$ with local coordinate $u$ as in Theorem E.36. Then
@@
This follows from Proposition E.37 and the Cauchy–Schwartz inequality. ∎
@@
The following is then a consequence of Proposition F.4 and the bound in Proposition F.7.
Corollary F.9.
@@
F.3 Soft Localization to an Open Set
For our next simplification we’ll deal with the difficulty of directly imposing the restriction of samples to an open neighborhood $U$ of $w^{*}$. We’ll deal this in practice by judicious choice of the prior $\varphi$ on $W$, following [Lau et al., 2025, Section 4.2]. Choose an embedding $\iota\colon W\hookrightarrow\mathbb{R}^{D}$ for some natural number $D$ and define $\varphi_{\sigma^{2}}$ by restricting a $D$-dimensional Gaussian centered at $\iota(w^{*})$ with variance $\sigma^{2}$ to $W$.
Remark F.10.
@@ F.3 Soft Localization to an Open Set
Since $\varphi_{\sigma^{2}}$ is still supported everywhere on $W$ Corollary F.9 still applies and the large $n\beta$ limiting behaviour of the estimator is independent of $\sigma^{2}$ in probability. The rate of convergence, however, depends on the value of $\sigma^{2}$: for any open neighborhood $U$ of $w^{*}$, any $n\beta,T$ and any $\varepsilon>0$ we can choose $\sigma^{2}$ small enough that the samples $y_{1},\ldots,y_{T}$ are all contained in $U$ with probability $1-\varepsilon$.
F.4 Estimators from Stochastic Gradient Langevin Dynamics
We will now discuss the LLC estimators that are practically computable, corresponding to the estimator defined in [Lau et al., 2025, Section 4.4]. These estimators are defined using stochastic gradient Langevin dynamics (SGLD) Welling and Teh [2011] as a substitute for sampling from the generalized posterior associated to the Gaussian prior $\varphi_{\sigma^{2}}$.
@@ F.4 Estimators from Stochastic Gradient Langevin Dynamics
    SGLD with constant step size is known, even under strong convexity assumptions on the objective function, to have limiting behaviour that differs from that of Langevin flow (as shown, for instance, in Brosse et al. [2018]). Therefore the SGLD based estimator is likely to be asymptotically biased with bias depending on the step size $\varepsilon$. Can this bias be quantified?
@@
    If one uses decaying step size then there are results implying almost sure convergence to the posterior distribution as the chain length $T\to\infty$ Teh et al. [2016]. These results, however, make assumptions that do not typically hold for the neural network models we are most interested in (see [Hitchcock and Hoogland, 2025, Appendix B] for a detailed discussion). What can one say upon relaxing these assumptions?
Remark F.13.
@@
If we use a subsequence $w_{B},\ldots,w_{T}$ for some initial point $B>0$ (the burn-in parameter) as a substitute for sampling from the posterior distribution $\exp(-n\beta G_{n}(w))\varphi_{\sigma^{*}}(w){\rm d}w$ on $W$ then we obtain the estimator that we compute in practice by analogy with the one defined in Definition F.11. For convenience we summarize the hyperparameters that this estimator depends on a table below.
@@
Appendix G Further LLC Estimate Data
@@
In this appendix we present our LLC estimates during training for models trained while varying the hyperparameters $\alpha$ and $\gamma$ and while keeping these hyperparameters fixed and varying the random seed. We present four ensembles of figures.
  • Figure 9 shows LLC estimates for independently sampled models with $\gamma=0.975$ and $\alpha=0.68$, with LLC estimation performed on distribution, i.e. with the same hyperparameters.
  • Figure 10 shows LLC estimates for models where estimation was performed with $\gamma=0.98$ and $\alpha=0.5$.
  • Figure 11 shows LLC estimates for models where estimation was performed with $\gamma=0.98$ and $\alpha=0$.
  • Figure 12 compares the LLC estimates with $\gamma=0.98$ and $\alpha=0.5$ to the weight norms of the model. We see that the weight norms are distinct from the LLC estimates. In particular they always increase monotonically during training and do not reflect the phase transitions to as significant an extent as the LLC estimates.
Figure 9: On-distribution LLC (blue line) estimated with $\gamma=0.975$ and $\alpha=0.68$ for runs with different seeds. The brown curve shows the regret, and the blue, beige and pink regions indicate phase 1, 2b and 3 respectively.
Figure 10: LLC (blue line) estimated with $\gamma=0.98$ and $\alpha=0.5$ for models trained with values of $\alpha$ and $\gamma$ indicated in the titles of the panels. The runs that share $\alpha$ and $\gamma$ differ only by the random seed with which they were trained. The brown curve shows the regret, and the blue, beige and pink regions indicate phase 1, 2b and 3 respectively.
Figure 11: LLC (blue line) estimated with $\gamma=0.98$ and $\alpha=0$ for models trained with values of $\alpha$ and $\gamma$ indicated in the titles of the panels. The runs that share $\alpha$ and $\gamma$ differ only by the random seed with which they were trained. The brown curve shows the regret, and the blue, beige and pink regions indicate phase 1, 2b and 3 respectively.
@@
Figure 12: LLC estimates (blue curves) with $\gamma=0.98$ and $\alpha=0.5$ for models trained with values of $\alpha$ and $\gamma$ indicated in the titles of the panels. The runs that share $\alpha$ and $\gamma$ differ only by the random seed with which they were trained. The shaded areas indicate the phase the model is in, with blue, beige and pink indicating phases 1, 2b and 3 respectively. These plots also show the weight norms (orange curves), which monotonically increase throughout training.
Appendix H Nonlinear relationship between LLC estimates and regret
@@
Figure 13: Distribution of the random variable $d$ representing the difference in the rate of change of the LLC estimate with respect to the regret between phases 1 and 2b, and between phases 2b and 3, measured across 21 different iid runs for which $\gamma=0.975$ and $\alpha=0.68$.
@@
where $\hat{\lambda}_{i}$ and $G_{i}$ are the average LLC estimates and average regrets among checkpoints in phase $i$ respectively. Under the hypothesis that $G_{i}$ and $\hat{\lambda}_{i}$ are linearly related the random variable $d$ would have expected value zero for all $\alpha$ and $\gamma$ values. We may test this using the mean of $d$ across a set of independent samples with fixed $\alpha$ and $\gamma$. We have 21 runs with $\alpha=0.68$ and $\gamma=0.975$, and show the distribution of $d$ for llcs computed on-distribution in Figure 13. Since these samples are iid we may perform a one sample t-test and reject the hypothesis that $\mathbb{E}(d)=0$ with significance $p=0.00005$.
Appendix I Sensitivity of automatic phase detection
@@
Figure 14: A plot showing the average number of checkpoints between Phase 1 and Phase 2b (blue), and the average number of checkpoints between Phase 2b to Phase 3 (orange). Models were trained with hyperparameters $\alpha=0.68,\gamma=0.975$. Training runs are only included if the policy was detected to be in all of phase 1, phase 2b and phase 3 at some point during training.
From Figure 14 we can see that for $0.10\leq\delta\leq 0.16$ the number of checkpoints between phase transitions is roughly constant (recall that spacing between checkpoints is logarithmic). In particular our LLC estimates within each phase are not significantly sensitive to the value of the threshold $\delta$ within this range.
When $\delta$ exceeds 0.16, we spuriously detect phase 2b early – our threshold has exceeded the distance between the two phases. Since we only include training runs where all three phases are present, for values of $\delta$ sufficiently small the sample size begins to decrease as fewer samples exceed the threshold for detecting the middle phase at some point during training; this is why the plots are not monotonic. Below $\delta=0.05$, no phases are detected in any of the runs, and the average is undefined.

### v2

@@ E.1 Asymptotically Unbiased Estimators
This follows from Theorem D.43. Indeed, by the weak law of large numbers
@@
where the expectation value is taken with respect to the generalized tempered posterior to $U$: $y_{j}\sim\Omega_{n,\beta}|_{U}$. Theorem D.43 then tells us that this expected value converges to $\lambda(U)$ in probability as $n\to\infty$, as required. ∎
E.2 The Annealed Posterior
Let’s discuss a variant of the WBIC-based LLC estimator with the same limiting properties as we saw in Proposition E.4. We will replace the tempered posterior in our definition with the annealed posterior (see Seung et al., 1992 for a discussion of this concept from its origin in statistical mechanics).
Definition E.5.
@@ E.2 The Annealed Posterior
Proposition E.7.
Fix a log resolution and a chart $U$ with local coordinate $u$ as in Theorem D.36. Then
@@
This follows from Proposition D.37 and the Cauchy–Schwartz inequality. ∎
@@
The following is then a consequence of Proposition E.4 and the bound in Proposition E.7.
Corollary E.9.
@@
E.3 Soft Localization to an Open Set
For our next simplification we’ll deal with the difficulty of directly imposing the restriction of samples to an open neighborhood $U$ of $w^{*}$. We’ll deal this in practice by judicious choice of the prior $\varphi$ on $W$, following Lau et al. [2025, §4.2]. Choose an embedding $\iota\colon W\hookrightarrow\mathbb{R}^{D}$ for some natural number $D$ and define $\varphi_{\sigma^{2}}$ by restricting a $D$-dimensional Gaussian centered at $\iota(w^{*})$ with variance $\sigma^{2}$ to $W$.
Remark E.10.
@@ E.3 Soft Localization to an Open Set
Since $\varphi_{\sigma^{2}}$ is still supported everywhere on $W$ Corollary E.9 still applies and the large $n\beta$ limiting behaviour of the estimator is independent of $\sigma^{2}$ in probability. The rate of convergence, however, depends on the value of $\sigma^{2}$: for any open neighborhood $U$ of $w^{*}$, any $n\beta,T$ and any $\varepsilon>0$ we can choose $\sigma^{2}$ small enough that the samples $y_{1},\ldots,y_{T}$ are all contained in $U$ with probability $1-\varepsilon$.
E.4 Estimators from Stochastic Gradient Langevin Dynamics
We will now discuss the LLC estimators that are practically computable, corresponding to the estimator defined in Lau et al. [2025, §4.4]. These estimators are defined using stochastic gradient Langevin dynamics [SGLD; Welling and Teh, 2011] as a substitute for sampling from the generalized posterior associated to the Gaussian prior $\varphi_{\sigma^{2}}$.
@@ E.4 Estimators from Stochastic Gradient Langevin Dynamics
    SGLD with constant step size is known, even under strong convexity assumptions on the objective function, to have limiting behaviour that differs from that of Langevin flow [as shown, for instance, in Brosse et al., 2018]. Therefore the SGLD based estimator is likely to be asymptotically biased with bias depending on the step size $\varepsilon$. Can this bias be quantified?
@@
    If one uses decaying step size then there are results implying almost sure convergence to the posterior distribution as the chain length $T\to\infty$ [Teh et al., 2016]. These results, however, make assumptions that do not typically hold for the neural network models we are most interested in [for a detailed discussion, see Hitchcock and Hoogland, 2025, Apx. B]. What can one say upon relaxing these assumptions?
Remark E.13.
@@
If we use a subsequence $w_{B},\ldots,w_{T}$ for some initial point $B>0$ (the burn-in parameter) as a substitute for sampling from the posterior distribution $\exp(-n\beta G_{n}(w))\varphi_{\sigma^{*}}(w){\rm d}w$ on $W$ then we obtain the estimator that we compute in practice by analogy with the one defined in Definition E.11. For convenience we summarize the hyperparameters that this estimator depends on a table below.
@@
Appendix F Further LLC Estimate Data
@@
In this appendix we present our LLC estimates during training for all 25 independent seeds with $\alpha=0.6,\gamma=0.98$. We present two ensembles of figures.
  • Figure 7 shows LLC estimates for 25 independently sampled models. We also show the behavior of the weight norms through training. We observe that the weight norms increase monotonically as the model trains and do not exhibit the ”staircase” behaviour of the LLC estimates, nor the decline within phases that we typically observe for LLC estimation.
  • Figure 8 shows LLC estimates for models where estimation was performed with a distinct value of $\alpha$ to that used in training (as described in Remark 6.2), specifically with $\alpha=0$. In other words during LLC estimation we only consider the behavior of the model in states where the goal is in the top-left corner. Note that in such states we cannot distinguish between phases 2 and 3 by means of the regret anymore: these phases are both optimal and receive zero regret.
The shaded areas indicate phases, whose detection was described in Section 6.4.1. The blue, beige and pink shaded regions indicate phases 1, 2b and 3 respectively.
We also show in green an additional phase that appears infrequently. Phase 2c exhibits optimal behaviour exactly in the half space where the mouse $x$-position in greater than or equal to the cheese $x$-position (or, alternatively, the identical condition with the $y$-coordinate). For the purpose of phase detection this is characterized by the region in policy space where $\pi(R)=0$ in all states, $\pi(D)=0$ if the mouse is not to the left of the cheese and not above the cheese, and $\pi(U)=0$ if the mouse is not to the left of the cheese and not below the cheese.
Remark F.1.
Observe that when estimating LLCs off distribution the transition between phase 2 and phase 3 is frequently visible through a spike in the LLC estimate, or a rapid change in the derivative, even though the policy does not distinguish between these two phases. The LLC estimates within the two phases however exhibit much higher variance both within and between seeds than we saw in the case of on-distribution LLC estimation. It may be possible to ameliorate this through more judicious choice of hyperparameters but doing so is not within the scope of the present work.
Remark F.2 (Environment complexity).
By enhancing our toy model slightly one could explore the ability of the LLC estimator to detect out-of-distribution generalization, meaning here performance in classes of states not seen during training. For a simple example consider the extension of the model presented here to maze environments (as in the original cheese-in-the-corner example presented in Abdel Sadek et al., 2025). One can allow the distribution of initial states used during training and LLC estimation to differ, but now changes in the LLC may detect changes in performance in out-of-distribution states.
@@
Figure 7: LLC estimates (blue curves) for models with $\gamma=0.98$ and $\alpha=0.6$.These plots also show the weight $L^{2}$ norms (orange curves), which monotonically increase throughout training, and the regret curves in pink. The weight norms range from 70 to $7$.
Figure 8: LLC (blue line) estimated with $\gamma=0.98$ and $\alpha=0$ for models trained with values of $\alpha$ and $\gamma$ indicated in the titles of the panels. The runs that share $\alpha$ and $\gamma$ differ only by the random seed with which they were trained. The brown curve shows the on-distribution regret and the green curve shows the regret measured with $\alpha=0$.
Appendix G Nonlinear relationship between LLC estimates and regret
@@
Figure 9: Distribution of the random variable $d$ representing the difference in the rate of change of the LLC estimate with respect to the regret between phases 1 and 2b, and between phases 2b and 3, measured across 25 different iid runs for which $\gamma=0.98$ and $\alpha=0.6$.
@@
where $\hat{\lambda}_{i}$ and $G_{i}$ are the average LLC estimates and average regrets among checkpoints in phase $i$ respectively. Under the hypothesis that $G_{i}$ and $\hat{\lambda}_{i}$ are linearly related the random variable $d$ would have expected value zero for all $\alpha$ and $\gamma$ values. We may test this using the mean of $d$ across a set of independent samples with fixed $\alpha$ and $\gamma$. We have 25 runs with $\alpha=0.6$ and $\gamma=0.98$ and show the distribution of $d$ for llcs computed on-distribution in Figure 9. Since these samples are iid we may perform a one sample t-test and reject the hypothesis that $\mathbb{E}(d)=0$ with significance $p=1.39\times 10^{-4}$.
Appendix H Sensitivity of automatic phase detection
@@
Figure 10: A plot showing the median number of checkpoints between Phase 1 and Phase 2b (blue), and the median number of checkpoints between Phase 2b to Phase 3 (orange). Models were trained with hyperparameters $\alpha=0.6,\gamma=0.98$. Training runs are only included if the policy was detected to be in all of phase 1, phase 2b and phase 3 at some point during training.
From Figure 10 we can see that for $0.08\leq\delta\leq 0.16$ the number of checkpoints between phase transitions is roughly constant (recall that spacing between checkpoints is logarithmic). In particular our LLC estimates within each phase are not significantly sensitive to the value of the threshold $\delta$ within this range.
When $\delta$ exceeds 0.16 we spuriously detect phase 2b early – our threshold has exceeded the distance between the two phases. Since we only include training runs where all three phases are present, for values of $\delta$ sufficiently small the sample size begins to decrease as fewer samples exceed the threshold for detecting the middle phase at some point during training; this is why the plots are not monotonic. Below $\delta=0.05$ no phases are detected in any of the runs, and the average is undefined.